OpenStudy (anonymous):

Which is the equation of a circle with the center at (4, 7) and the radius 4 units long? (x + 4)2 + (y + 7)2 = 16 (x − 4)2 + (y − 7)2 = 16 (x + 4)2 + (y + 7)2 = 4 (x − 4)2 + (y − 7)2 = 4 HELP PLEAS!!!

OpenStudy (anonymous):

I believe it is the second one: (x − 4)2 + (y − 7)2 = 16 We know that the center is 4,7, so it will be x-4 and y-7 C=4, so C^2=16

OpenStudy (anonymous):

Thanks I actually had that answer but I wanted to check with others.